OpenStudy (keana):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5). a. y = 5x − 15 b. y = 5x + 15 c. y = 1/5x + 21/5 d. y = 1/5x − 29/5

OpenStudy (raden):

parallel means has the same slope. write that equation is y = 5x + c now distribute the point (4,5), get 5 = 5(4) + c 5 = 20 + c c = 5 - 20 c = -15 therefore, this line is y = 5x - 15 does that make sense ?

OpenStudy (keana):

now i get it. i wasn't distributing properly. thanks!

OpenStudy (raden):

wlcm